Which charge could not be found on an object?
(1) -0.8 x 10-19 C (3) +1.6 x 10-19 C
(2) -1.6 x 10-17 C (4) +3.2 x 10-17 C
Proof: Explain your reasoning/show calculations.
1
Expert's answer
2020-04-20T10:40:37-0400

The elementary charge (the quanta of charge) equals "e=1.6\times 10^{-19}\:\rm C."

Hence, total charge of an object

"q=Ne,"

here "N" is integer.

Let's check:

(1)

"N=\frac{0.8\times 10^{-19}}{1.6\times 10^{-19}}=1/2 \; (\rm impossible)"

(2)

"N=\frac{1.6\times 10^{-17}}{1.6\times 10^{-19}}=100 \; (\rm possible)"

(3)

"N=\frac{1.6\times 10^{-19}}{1.6\times 10^{-19}}=1 \; (\rm possible)"

(4)

"N=\frac{3.2\times 10^{-17}}{1.6\times 10^{-19}}=200 \; (\rm possible)"